# LLM-Text Watermarking based on Lagrange Interpolation

### v1

@@ 5.1 Maximum Collinear Points Problem
The adversarial model is illustrated in Figure 1. In this model, the adversary has full control over the text and may apply arbitrary modifications, including insertions and deletions. However, extensive editing is often counterproductive, as it effectively amounts to rewriting the entire text from scratch. In practice, it is reasonable to assume that the adversary will aim to preserve as much of the original content as possible to minimise effort – after all, the motivation behind plagiarism is typically to save time by reusing someone else’s work. Interestingly, if the original watermark is available, the secret $K$ can be recovered directly using Lagrange interpolation. However, a verifier who only knows the key $s_{k}$ faces a more challenging task, as the extracted watermark may have been tampered with by an adversary. Effectively, the verifier deals with an instance of the MCP problem.
@@
Let us consider the Hashing-based algorithm (see Algorithm LABEL:alg:hash_MCP). It determines the largest subset of collinear points among $N$ given points in a 2D plane. The algorithm iterates through each point, treating it as a reference. For every other point, it computes the slope relative to the reference and stores the slope count in a hash map. This hash map allows for quick lookups and updates, ensuring that collinear points are efficiently counted. The maximum count of any slope within the hash map, plus the reference point itself, gives the largest number of collinear points for that reference. By iterating through all points, the algorithm determines the global maximum. Note that vertical lines can be ignored as they are not parts of secret sharing.
@@ 6 Security Evaluation
6.1 Algebraic Analysis
By now, the reader should notice that the proposed watermarking scheme relies on two linear components: the LFSR and the function $f(x)$. The only nonlinear element arises from the division of the token set into ${\cal V}_{0}$ and ${\cal V}_{1}$, which occurs when cryptographic hashing or encryption is employed, as discussed in Section 4.3. It is important to observe that Algorithm 5 uses fixed sets ${\cal V}_{0}$ and ${\cal V}_{1}$ throughout the entire embedding process. This design choice introduces a potential vulnerability, as it enables the adversary to mount the following attack.
Algebraic Attack on Watermarking Scheme
  1. 1.
    Target the most frequent tokens that appear in the attacked LLM-generated text. Determine $n$ positions, where the top most frequent tokens occur.
  2. 2.
    For each position and its token, write a binary linear relation of the form:
    $\bigoplus_{i=1}^{n}\delta_{i}s_{i,k}=w_{j},$
    where $s_{k}=(s_{1,k},\ldots,s_{n,k})$ is the LFSR seed, and $w_{j}$ is an unknown bit. The value $w_{j}$ is either 00 or $1$, depending on whether the token belongs to ${\cal V}_{0}$ or ${\cal V}_{1}$.
  3. 3.
    Collect the resulting system of linear equations and solve them for guessed values of $w_{j}$. If the adversary observes the same token in $n$ positions, they can solve the system for both $w_{j}=0$ and $w_{j}=1$, and be confident that one solution is correct.
  4. 4.
    Verify the candidate solution by substituting the computed $\tilde{s}_{k}$ into the LFSR and checking whether the generated sequence matches the observed token positions. This reveals how watermarking bits are assigned to tokens, assuming the LFSR governs watermark generation.
  5. 5.
    Using the identified tokens, assemble $n$ binary relations associated with the function $f(x)$ and recover the secret identity $K$. This is feasible when some tokens carrying values $f(x_{i})$ have been identified via their membership in ${\cal V}_{0}$ or ${\cal V}_{1}$. As a result, the watermarking scheme can be fully compromised.
Algebraic attacks can be mitigated using simple yet effective countermeasures, as outlined below.
  • $\bullet$
    Instead of using an LFSR, it is advisable to employ a cryptographically secure NFSR, such as those proposed in the CAESAR competition [3]. While an adversary may still attempt to construct algebraic relations for the most frequent tokens, each guess of watermark bits leads to highly nonlinear equations, rendering the system intractable. For security-critical applications, it is recommended to use provably secure cryptographic pseudorandom bit generators.
  • $\bullet$
    As an alternative – or in addition – to the above, the fixed sets ${\cal V}_{0}$ and ${\cal V}_{1}$ can be reshuffled periodically. For instance, the token set can be repartitioned every $n$-bit or $2n$-bit block. Furthermore, different partitions can be employed for the $x_{i}$ blocks generated by the LFSR and for the $f(x_{i})$ blocks. In the most extreme approach, ${\cal V}_{0}$ and ${\cal V}_{1}$ can be reshuffled for every individual token. However, this significantly increases computational overhead, especially during watermark extraction.
Watermarking Scheme Immune to Algebraic Attacks Given the watermarking scheme with embedding and extraction described in Algorithms 5 and 5, the following modifications make the scheme immune to algebraic attacks: 1. Apply a nonlinear feedback shift register (NFSR) instead of an LFSR to prevent recovery of the seed $s_{k}$. 2. Replace fixed token space partitions with dynamic ones, depending on the type of token: $\bullet$ For tokens corresponding to $x_{i}$ generated by NFSR, use a fixed partition of ${\cal V}$. For instance, assign a token $v$ with identity $ID_{v}$ a watermark bit $m=\bigoplus_{j=1}^{n}h_{j}$, where $(h_{1},\ldots,h_{n})=\mathrm{SHA3}(s_{k},ID_{v})$. $\bullet$ For $y_{i}=f(x_{i})$, create a new partition for each $i$, breaking the linear dependencies induced by $f(x)$. For example, define the token partition using a hash $H(s_{k},x_{i})$. A specific instance might compute $(h_{1},\ldots,h_{n})=\mathrm{SHA3}(s_{k},x_{i},ID_{v})$. 3. This design enables efficient watermark extraction with only a slight increase in complexity. The verifier first applies the partitioning for tokens $x_{i}$ generated by NFSR. Once all $x_{i}$ are identified, the verifier extracts the watermark values $y_{i}$ from the corresponding $n$-token blocks.
6.2 Probability of Watermark Extraction
@@
Let us consider the extent to which our adversary must tamper with our watermarking to destroy the watermark. In other words, we aim to determine the threshold at which the watermark’s owner can still prove that the LLM-generated text has been watermarked by them or extract the owner’s identity $K$. We can extract $N=F+R$ points, where $F$ belong to the straight line $f(x)$ and $R$ have been tampered with by the adversary. It is reasonable to assume that the $R$ points are randomly scattered on the plane $GF(2^{n})\times GF(2^{n})$. Clearly, the $K$ extraction is successful if any subset of random points forming a straight line is no larger than $F$.
@@
7 Extensions and Variants
7.1 Simplified Watermark Extraction
Algorithm 5 recovers watermarks by identifying the values $\mathbb{L}=(x_{1},\ldots,x_{N})$, which serve as the $x$-coordinates of the points. After locating an $x_{i}$ that was generated by LFSR($s_{k}$), the subsequent block of $n$ bits is extracted and interpreted as $f(x)$. In other words, our MCP algorithm operates on at most $N$ points, as those with incorrect $x_{j}$ values are ignored. This selective approach is particularly effective when dealing with long texts. However, for shorter texts, we may be inclined to run the MCP algorithm directly – without filtering out points with invalid $x$-coordinates. The alternative idea is to consider $12n(N-1)+$ 1 points, which are generated by sliding a $2n$-bit window bit by bit across the entire $2nN$-bit watermark. Our experiments confirmed expectations: apart from the $N$ collinear original points, the additional points produced through bit shifts behave like random points and do not interfere with the collinear structure. However, special attention must be paid to small Galois fields such as $GF(2^{6})$. In such cases, the random selection of $x$-coordinates may result in duplicate points, and the quality of randomness from the sliding-window operation may be questionable. Table 3 shows results, when $x$-coordinates are selected (pseudo)-randomly.
Arithmetics # Collinear Points # Points by Shifting
5 49
$GF(2^{6})$ 10 109
20 (18) 229
$GF(2^{12})$ 10 109
80 1897
$GF(2^{32})$ 40 2497
100 6337
Table 3: Collinear points when the MCP algorithm is fed with points generated by shifting
Note that the red entry for $GF(2^{6})$ indicates that our MCP algorithm returned 18, which is clearly incorrect. Since we used pseudorandom generation, there is a high likelihood that some of the 20 random points are duplicates, given that there are only 64 possible $x$-coordinates to choose from.
7.2 Double-Resolution Watermarking
We embed watermarks by defining a sequence of points $(x_{i},f(x_{i}))\in GF^{2}(2^{n})$; $i=1,\ldots,N$. When using our simplified extraction method, the MCP algorithm must process $12n(N-1)+$ 1 points. However, the number of points generated by shifting a $2n$-bit window can be reduced to $n(N-1)+1$ by overlapping point coordinates. In other words, instead of using the coordinates $x_{1},x_{2},\ldots,x_{N}$ generated by an LFSR, we can let $f(x)$ recursively generate the coordinates as $x,f(x),f^{2}(x),\ldots,f^{N-1}(x)$. Consequently, the sequence of points becomes
($\underset{P_{1}}{\underbrace{(x,f(x))}},\ \underset{P_{2}}{\underbrace{(f(x),f^{2}(x))}},\ \ldots,\ \underset{P_{N}}{\underbrace{(f^{N-1}(x),f^{N}(x))}}$
The sequence can be compressed into
$(x,f(x),f^{2}(x),\ldots,f^{N-1}(x),f^{N}(x))$
It still preserves the $N$ collinear points, while reducing the number of “random” points introduced by the shifting operation. There are, however, some caveats to consider. First, repeated application of the function $f(x)$ leads to a cycle. If this cycle is shorter than $N$, then a new initial $x$ must be selected. Second, and more importantly, this design may have security implications that must be carefully addressed.
@@ 8 Experiments
Given a short text, such as a paragraph consisting of 300 tokens, we can embed a 300-bit watermark. In practice, a key challenge lies in selecting the appropriate dimension of the Galois field used for the computation. If we choose $GF(2^{6})$, each point occupies 12 bits. This means the text can accommodate 25 watermark points. According to Table LABEL:tab:mcp, only 5 of these points need to survive for the correct watermark line to be recovered, while the remaining 20 points may be replaced by fake or random ones without affecting recoverability. If we instead use $GF(2^{8})$, the total number of watermark points decreases to 18. In this case, as shown in Table LABEL:tab:mcp, just 4 genuine points are enough to reconstruct the correct line, while the remaining 14 random points form lines with a maximum of only 3 points. With $GF(2^{12})$, the number of embedable points drops further to 12. Still, only 3 genuine watermark points are needed to reconstruct the original line, as 9 random points only form trivial two-point lines. A broader discussion considering different text lengths – such as a paragraph, a page, and a 10-page report – is provided in Table 5. It is evident that for longer documents, using Galois fields of higher dimension is advantageous, as the appearance of a three-point line becomes strong evidence of the watermark’s presence and authenticity.
@@
$GF(2^{6})$ 25 5 20
Paragraph 300 $GF(2^{8})$ 18 4 14
$GF(2^{12})$ 12 3 9
$GF(2^{6})$ 58 7 51
Page 700 $GF(2^{8})$ 43 5 38
$GF(2^{12})$ 29 3 26
$GF(2^{6})$ 583 20 563
Report 7,000 $GF(2^{8})$ 437 5 432
(10 pages) $GF(2^{12})$ 291 4 287
$GF(2^{16})$ 218 3 215
@@
9 Muli-Secret Watermarking
@@ 9.1 Multiple Lines
Using a single line $f(x)=ax+b\in GF(2^{n})$, it is possible to obtain the secret $K=(a\|b)\in\{0,1\}^{2n}$. Further expansion of the secret is possible by embedding $t$ distinct lines instead of just one. In this case, our MCP algorithm requires a minor adjustment to track the top $t$ straight lines that yield the highest point counts. This results in a collection of lines $\{f_{i}(x)=a_{i}x+b_{i}\;|\;i=1,\ldots,t\}$ and their corresponding secrets $\{K_{i}=(a_{i}\|b_{i})\;|\;i=1,\ldots,t\}$. To form a single $2tn$-bit secret, we must determine an appropriate order for concatenation. A simple method proceeds as follows:
@@
    For each secret $K_{i}$, where $i=1,\ldots,t$, compute its digest: $d_{i}=H(K_{i})$, where $H()$ is a cryptographic hashing.
@@
    Sort the digests $\{d_{1},\ldots,d_{t}\}$ in increasing order and reorder the corresponding secrets $K_{i}$ accordingly.
@@
    Concatenate the reordered secrets into a single sequence.
@@ 9.2 High Degree Polynomials
Watermarks are points $\{(x_{i},f(x_{i})\;|\;i=1,\ldots,N\}$. Clearly, to determine the polynomial $f(x)$, it is enough to know $t$ points. We can apply the well-know Lagrange Interpolation to reconstruct $f(x)$. The secret $K=(a_{0}\;\|\;a_{1}\;\|\;\ldots\;\|\;a_{t-1}\}$. The problem of recovery, however, becomes interesting when some points are corrupted by our adversary. The verifier in this case faces a task to identify a polynomial of degree $(t-1)$, which contains the maximum number of points (assuming that the adversary corrupted a relatively small portion of points). This is to say that we face the Maximum Co-Polynomial Points (MCPP) problem.
@@ 10 Conclusions
This work introduces a novel watermarking scheme for LLM-generated text that enables the recovery of a secret author identity through geometric encoding. The core idea is to represent the secret as a straight line $f(x)$ over a finite field, and embed a sequence of points $(x_{i},f(x_{i}))$ lying on this line into the LLM text. These points serve as the watermark, recoverable even after partial editing or adversarial manipulation. We demonstrate that the scheme is lightweight, easy to implement, and well-suited for applications with moderate security requirements. For security-critical scenarios, we enhance the robustness of the scheme by generating the $x_{i}$ coordinates using cryptographically secure pseudorandom number generators, and by introducing a variable partitioning of token sets for each $n$-token block. These modifications increase resistance to algebraic and probabilistic attacks.
The scheme is naturally extendable to support multiple secrets. One straightforward approach is to embed multiple lines, each corresponding to an individual secret. While effective, this method increases the probability that an adversary may generate a fake point that aligns with one of the embedded lines, potentially resulting in false positives. Moreover, if the goal is to produce a single unified secret, a method for determining the correct order of concatenation among the individual secrets must be established. To address this, we propose the use of higher-degree polynomials instead of lines. In this approach, the secret is encoded as the coefficients of a polynomial $f(x)$ and watermark points are sampled from this polynomial. This not only reduces the adversary’s chances of fabricating a point that lies on the curve—since a random point lies on a polynomial of the degree $t$ with probability only $\nicefrac{{1}}{{2^{n}}}$ – but also enables straightforward concatenation of the coefficients into a single, longer secret. The trade-off, however, is an increase in computational complexity: reconstructing a polynomial of the degree $t$ requires solving a more resource-intensive interpolation problem, especially as $t$ or the field size grows.
Overall, the proposed watermarking framework offers a flexible balance between efficiency and security, and opens the door to new watermarking paradigms based on algebraic structures. Future work may focus on optimising multi-secret recovery, enhancing resilience against editing attacks, and exploring hybrid schemes combining lines and polynomials to balance robustness and performance.

### v2

@@ 5.1 Maximum Collinear Points Problem
The adversarial model is illustrated in Figure 1. In this model, the adversary has full control over the text and may apply arbitrary modifications, including insertions and deletions. However, extensive editing is often counterproductive, as it effectively amounts to rewriting the entire text from scratch. In practice, it is reasonable to assume that the adversary will aim to preserve as much of the original content as possible to minimise effort – after all, the motivation behind plagiarism is typically to save time by reusing someone else’s work. Interestingly, if the original watermark is available, the identity $K$ can be recovered directly using Lagrange interpolation. However, a verifier who only knows the key $s_{k}$ faces a more challenging task, as the extracted watermark may have been tampered with by an adversary. Effectively, the verifier deals with an instance of the MCP problem.
@@
Let us consider the Hashing-based algorithm (see Algorithm 5.1). It determines the largest subset of collinear points among $N$ given points in a 2D plane. The algorithm iterates through each point, treating it as a reference. For every other point, it computes the slope relative to the reference and stores the slope count in a hash map. This hash map allows for quick lookups and updates, ensuring that collinear points are efficiently counted. The maximum count of any slope within the hash map, plus the reference point itself, gives the largest number of collinear points for that reference. By iterating through all points, the algorithm determines the global maximum. Note that vertical lines can be ignored as they are not parts of secret sharing.
@@ 6 Security Evaluation
According to our security model (see Figure 1), the adversary has full control over the LLM-generated text. They may freely modify, delete, edit, replace, or duplicate any part of it. In this setting, unconditional security is unattainable, as the adversary can always rewrite the content from scratch using the original as a guide. A more realistic goal is to ensure that an adversary – seeking to save time by reusing parts of the original text – is likely to leave behind fragments of the watermark that, with high probability, still allow recovery of the original author’s identity. Therefore, the primary security objective is to preserve authorship attribution. Indirectly, this means (1) maximizing the probability of recovering the embedded identity, and (2) increasing the effort required to fully erase authorship to a level comparable with rewriting the entire text from scratch.
@@
Let us consider the extent to which an adversary must tamper with the watermarked text to successfully destroy the watermark. In other words, we aim to determine the threshold beyond which the original owner can no longer prove authorship or extract their identity $K$ from the LLM-generated text. During extraction, we obtain a total of $N=F+R$, where $F$ are valid points that lie on the original straight line $f(x)$ and $R$ are spurious points resulting from two sources: (1) noise introduced by the extraction process itself, and (2) modifications introduced by the adversary. It is reasonable to assume that the $R$ points are randomly scattered on the plane $GF(2^{n})\times GF(2^{n})$. Clearly, the $K$ extraction is successful if any subset of random points forming a straight line is no larger than $F$.
@@ 7 Experiments
Given a short text, such as a paragraph consisting of 300 tokens, we can embed a 300-bit watermark. In practice, a key challenge lies in selecting the appropriate dimension of the Galois field used for the computation. If we choose $GF(2^{6})$, each point occupies 6 bits. This means the text can accommodate 50 watermark points. According to Table 4, only 6 of these points need to survive for the correct watermark line to be recovered, while the remaining 44 points may be replaced by fake or random ones without affecting recoverability. If we instead use $GF(2^{8})$, the total number of watermark points decreases to 36. In this case, as shown in Table 4, just 4 genuine points are enough to reconstruct the correct line, while the remaining 32 random points form lines with a maximum of only 3 points. With $GF(2^{12})$, the number of embedable points drops further to 25. Still, only 3 genuine watermark points are needed to reconstruct the original line, as 22 random points only form trivial two-point lines. A broader discussion considering different text lengths – such as a paragraph, a page, and a 10-page report – is provided in Table 4. It is evident that for longer documents, using Galois fields of higher dimension is advantageous, as the appearance of a three-point line becomes strong evidence of the watermark’s presence and authenticity.
@@
$GF(2^{6})$ 50 6 44
Paragraph 300 $GF(2^{8})$ 36 4 32
$GF(2^{12})$ 25 3 22
$GF(2^{6})$ 116 7 109
Page 700 $GF(2^{8})$ 87 5 82
$GF(2^{12})$ 58 4 54
$GF(2^{6})$ 1166 26 1140
Report 7,000 $GF(2^{8})$ 875 10 865
(10 pages) $GF(2^{12})$ 583 5 578
$GF(2^{16})$ 437 3 433
@@
8 Long Multi-bit Watermarking
@@ 8.1 Multiple Lines
Using a single line $f(x)=ax+b\in GF(2^{n})$, it is possible to obtain the identity $K=(a\|b)\in\{0,1\}^{2n}$. Further expansion of its length is possible by embedding $t$ distinct lines instead of just one. In this case, our MCP algorithm requires a minor adjustment to track the top $t$ straight lines that yield the highest point counts. This results in a collection of lines $\{f_{i}(x)=a_{i}x+b_{i}\;|\;i=1,\ldots,t\}$ and their corresponding identities $\{K_{i}=(a_{i}\|b_{i})\;|\;i=1,\ldots,t\}$. To form a single $2tn$-bit identity, we must determine an appropriate order for concatenation. A simple method proceeds as follows:
@@
    For each identity $K_{i}$, where $i=1,\ldots,t$, compute its digest: $d_{i}=H(K_{i})$, where $H()$ is a cryptographic hashing.
@@
    Sort the digests $\{d_{1},\ldots,d_{t}\}$ in increasing order and reorder the corresponding identitys $K_{i}$ accordingly.
@@
    Concatenate the reordered identitys into a single sequence.
@@ 8.2 High Degree Polynomials
Points $\{(x_{i},f(x_{i})\;|\;i=1,\ldots,N\}$ are translated into appropriate watermarking bits. Clearly, to determine the polynomial $f(x)$, it is enough to know $t$ points. We can apply the well-know Lagrange Interpolation to reconstruct $f(x)$. The identity $K=(a_{0}\;\|\;a_{1}\;\|\;\ldots\;\|\;a_{t-1}\}$. The problem of recovery, however, becomes interesting when some points are corrupted by our adversary. The verifier in this case faces a task to identify a polynomial of degree $(t-1)$, which contains the maximum number of points. This is to say that we face the Maximum Co-Polynomial Points (MCPP) problem.
@@ 9 Conclusions
This work introduces a novel watermarking scheme for LLM-generated text that enables the recovery of a identity author identity through geometric encoding. The core idea is to represent the identity as a straight line $f(x)$ over a finite field, and embed a sequence of points $(x_{i},f(x_{i}))$ lying on this line into the LLM text. These points serve as the watermark, recoverable even after partial editing or adversarial manipulation. We demonstrate that the scheme is lightweight, straightforward to implement, and well-suited for real-world applications.
The scheme is naturally extendable to support multiple identities. One straightforward approach is to embed multiple lines, each corresponding to an individual identity. While effective, this method increases the probability that an adversary may generate a fake point that aligns with one of the embedded lines, potentially resulting in false positives. Moreover, if the goal is to produce a single unified identity, a method for determining the correct order of concatenation among the individual identities must be established. To address this, we propose the use of higher-degree polynomials instead of lines. In this approach, the identity is encoded as the coefficients of a polynomial $f(x)$ and watermark points are sampled from this polynomial. This not only reduces the adversary’s chances of fabricating a point that lies on the curve—since a random point lies on a polynomial of the degree $t$ with probability only $\nicefrac{{1}}{{2^{n}}}$ – but also enables straightforward concatenation of the coefficients into a single, longer identity. The trade-off, however, is an increase in computational complexity: reconstructing a polynomial of the degree $t$ requires solving a more resource-intensive interpolation problem, especially as $t$ or the field size grows.
Overall, the proposed watermarking framework offers a flexible balance between efficiency and security, and opens the door to new watermarking paradigms based on algebraic structures. Future work may focus on optimising multi-identity recovery, enhancing resilience against editing attacks, and exploring hybrid schemes combining lines and polynomials to balance robustness and performance.